Mr.Lucero's cornbread recipe calls for 3 cups of cornmeal. He puts in 1 2/3 cups from one open bag and 2/3 cup from another open bag. How many cups of cormeal does Mr. Lucero need to add when he opens a third bag.
Answer by stanbon(48558)

Mr.Lucero's cornbread recipe calls for 3 cups of cornmeal. He puts in 1 2/3 cups from one open bag and 2/3 cup from another open bag. How many cups of cormeal does Mr. Lucero need to add when he opens a third bag.
-----------------------------------------
3 - (5/3 + 2/3) = 3 - 7/3 = 2/3 cups needed.
=========================================
Cheers,
Stan H.